# A Circuit-QED Lattice System with Flexible Connectivity and Gapped Flat Bands for Photon-Mediated Spin Models

### v1

@@ Appendix A Supplementary Transmission Data
Figure 8: Low-power half-wave transmission data. (a)-(b) Transmission data of the half-wave modes as a function of flux applied to Q1 while the remaining two qubits are detuned from the range of interest: (a) transmission magnitude near the degenerate flat bands and lowest-energy dispersive band and (b) transmission magnitude over the full set of half-wave modes. (c)-(d) Analogous data for flux applied to Q2. (e)-(f) Analogous data for flux applied to Q3. The mutiphoton-transition and transmon-ionization glitches seen in Fig. 6 are not present here due to the lower input power used ($5$ dB lower for the zoomed-in scans, $10$ dB lower for the wide scans), and the avoided crossings are smooth and continuous. Due to the intra-unit-cell wavefunctions in the different bands and the inequivalent locations of the three qubits, Q2 does not couple to the half-wave flat bands, and as a result, the large avoided crossing visible for Q1  and Q3  in (a) and (e), is not present when tuning Q2  in (c). Additionally, the mode at $4.95$ GHz is an example of a mode which has a node in one unit cell, and therefore does not couple to Q1 in (b), but does couple to Q3 in (f), which is located on the equivalent lattice site one unit cell over.
@@
Since large probe powers can lead to non-linear effects, a tradeoff must be made between visibility of small features and clean avoided crossings. At high powers, modes are clearly visible with few averages and low-transmission modes can be detected with higher confidence. However, when a qubit is swept through resonance with a mode the drive tone has the potential to ionize the qubit (see Appendix D) leading to glitches in the avoided crossings. At low powers transmon ionization is less likely but achieving equivalent SNR for small features is generally not possible, or extremely slow. The crossing between the half-wave mode at $\sim 4.96$ GHz and Q3 is a perfect example of this effect of measurement-tone power. In a higher-power scan, such as Fig. 6(c), the avoided crossing "breaks" when Q3 is close to resonance with the mode, but in a scan taken with $10$ dB lower drive power (Fig. 8(f)), the full crossing is cleanly visible.
@@
Numerical simulations of the photonic bound states, shown in Fig. 11 and discussed in Appendix C, indicate that Q3, and by similarity Q1, attain peak frequencies of roughly $9.85-9.9$ GHz, which are inside the full-wave band structure. As a result, modes in the highest dispersive band do not generally show full avoided crossings, merely deflections. Modes in this upper dispersive band are particularly difficult to observe directly in transmission, due in part to the broad parasitic modes situated within the band. For these modes, the phase response of the transmitted signal shown in Fig. 9(a)-(c)(iii) is stronger than the amplitude response that is typically used to characterize the other photonic modes.
Appendix B General Device Design
B.1 On-Site Energies
The resonator array featured in this work is composed of nine identical unit cells, each containing six CPW resonators. Each resonator is designed to have the first harmonic at $5$ GHz and the second at $10$ GHz, and unintended frequency differences between resonators would introduce disorder to the lattice which in turn can result in undesired localization. As such, the designs of the resonators utilized in our final device were intentionally chosen to minimize frequency disorder. More specifically, we use CPWs with wide center pins and gentle bend radii that are larger than the field standard to reduce sensitivity to disorder introduced during device fabrication [4]. Our CPWs have a center-pin width of $20$ $\mu$m, a gap width of $8.372$ $\mu$m, and bend radius of $90$ $\mu$m.
Physically realizing the quasi-1D target lattice on a square chip requires the usage of multiple resonator shapes throughout the device. The resonant frequency of a CPW is primarily set by the length of the resonator, and to leading order, the choice in resonator shape does not introduce frequency disorder. However, higher-order corrections to the resonant frequency can arise, e.g. due to the number of bends in the resonator and variations in the bend radii. To minimize the impact of frequency disorder between resonator shapes and the amount of calibration needed, we only use four distinct resonator shapes and we correct systematic frequency differences between them by adding or subtracting from the length of the CPW. In order to determine the corrections needed, we tested each resonator geometry individually. The CAD of the two test chips is shown in Fig. 10. In order to separate geometry-dependent frequency shifts from the effects of run-to-run variations in lithography exposure, both devices were exposed and etched on a single chip before being diced and separated. The half-wave frequency of the resonator with the fewest bends (Resonator 1 in Fig. 10(a)) differed from the other three designs by nearly $2\pi\times 20$ MHz: Resonator 1 had a frequency of $2\pi\times 4.8778$ GHz while Resonators 2-4 had frequencies of $2\pi\times 4.8957\pm 0.001$ GHz. For the final design, a length of $45.48\ \mu$m was removed from Resonator 1 to match its frequency with the rest of the resonators.
@@ B.1 On-Site Energies
Figure 10: Resonator Shape Test. (a)-(b) CAD of the two shape-test devices that were fabricated and measured to determine the frequencies of the four distinct CPW resonator shapes used. The CPW resonators, shown in red, are each capacitively coupled to an input and an output line, shown in gray. The couplers are shown in dark blue. Two capacitor pads, highlighted in purple, are coupled to the side of the resonator to account for the shift in the frequency of the resonator from adding a transmon. High-frequency $\lambda/4$ resonators, used to terminate unused coupler parts, are highlighted in yellow. Resonator $1$ is found to have a significantly lower frequency than the other three resonators. This effect is compensated for in the full resonator array by removing $45.48\ \mu$m from all resonators with the same shape as Resonator $1$.
Another potential source of disorder arises from the capacitance between neighboring resonators, since the coupling capacitors on the device, highlighted blue in Fig. 10, induce on-site energy shifts in the resonators. To keep this contribution uniform, all resonators must have the same number of coupling capacitors [58, 18], two at each end for the case of the device shown here. Therefore, the same 3-way coupling-capacitor geometry is used at the ends of all resonators, both for sites in the bulk which naturally have four neighbors, as well as for sites at the edge of the device, where some of the neighboring resonators are absent. Unused ports on the 3-way capacitors are either connected to input/output feedlines, or terminated with a high-frequency $\lambda/4$ resonator, which does not participate in the lattice due to the large detuning.
@@
B.2 Hopping Parameters
Following the arguments in Refs. [46, 58], the effective tight-binding hopping parameter for two capacitively coupled CPW resonators is given by:
$t_{\mu}=-\frac{1}{2}\omega_{\mu}C_{c}\tilde{\Phi}_{\mu}^{2},$ (18)
where $C_{c}$ is the coupling capacitance, and $\tilde{\Phi}_{\mu}$ denotes the value of the mode function on the two resonator ends participating in the coupling capacitor. To convert this into a simpler form in terms of fundamental properties of the CPW and the harmonics of the resonator, we evaluate $\tilde{\Phi}_{\mu}$ using the normalization condition
[…]
However, the form above now depends on the the capacitance per unit length $c_{\ell}$, which cannot be measured directly. A more convenient form can be obtained by eliminating $c_{\ell}$ and $L$ via their relationship to the impedance $Z_{0}$ of the CPW transmission line and the fundamental frequency $\omega_{1}$ of the resonator [64]: $L\sqrt{\ell_{\ell}c_{\ell}}=\pi/\omega_{1}$ and $Z_{0}=\sqrt{\ell_{\ell}/c_{\ell}}$, where $\ell_{\ell}$ is the inductance per unit length of the CPW. Hence, $2/c_{\ell}L=2\omega_{1}Z_{0}/\pi$, and
$t_{\mu}\approx-\frac{1}{\pi}\omega_{1}\omega_{\mu}C_{c}Z_{0}.$ (20)
To determine the predicted zeroth-order hopping rate in our device, we use the resonant frequencies determined from the shape-test devices (or finite element simulations in HFSS), $Z_{0}=50\ \Omega$, and a finite element simulation of the coupling capacitance using Ansys Maxwell. Our three-way coupler with a gap of $2\leavevmode\nobreak\ \mu m$ is simulated to have a $19.5$ fF capacitance between each pair of resonators, from which $|t_{1}^{(0)}|$ is predicted to be $\sim 2\pi\times 47$ MHz and $|t_{2}^{(0)}|$ is predicted to be $\sim 2\pi\times 94$ MHz. For the main device, the empirically determined hopping rates for the half-wave and full-wave modes are determined through comparison of the experimentally-measured photonic bands with the frequency-dependent hopping model of the target lattice (see Section V.2). These values of $|t_{1}^{(0)}|\approx 2\pi\times 40$ MHz and $|t_{2}^{(0)}|\approx 2\pi\times 82$ MHz are reasonably close to the simulated values predicted by circuit theory and finite-element simulation. The discrepancy may be due to slight over-exposure of the device during photolithography, or due to inaccuracy in the Ansys Maxwell capacitance simulation at the level of $\sim 1$ fF.
B.3 Transmon Design
Before the design of the transmon qubits embedded in the CPW lattice is discussed in detail, we first introduce some of the traditional parameters of a transmon, as defined in Refs. [56, 9, 65]. The frequency of the transition between the ground and first-excited state is approximately given by
@@ B.3 Transmon Design
The total capacitance takes into account not only the direct capacitance between the two paddles but also the indirect coupling through the center pin of the CPW resonator and the ground planes of the device [65]. For flux-tunable transmons, the effective total $E_{J}$ is dependent on the individual Josephson energies of the two junctions in the SQUID loop and the external applied flux. The full form for the effective $E_{J}$ of a tunable transmon is given by
@@
In order to suppress the effect of charge noise in the substrate, a transmon must operate in the regime where $E_{J}/E_{C}\gg 1$, typically $20$ or larger [56, 9]. This requires a large $C_{\Sigma}$, which can either be achieved using large-area capacitor paddles or a digitated capacitor. Our design uses a digitated capacitor in order to make the qubit compact and easier to incorporate in different resonator geometries. Conventional, non-digitated designs would need to be significantly wider, as much as $5-10$ times, to achieve an equivalent capacitance, which would severely restrict the ability to tightly pack resonators in the array [66]. For our design, the total capacitance between the paddles is determined to be $C_{\Sigma}=0.1174$ pF using Ansys Maxwell. Following from Eq. 22, $E_{C}$ is predicted to be $2\pi\times 165$ MHz. In a transmon, the two-photon transition from the ground state to the second-excited state is lower in frequency than the main transition by $\sim E_{C}/2$ [9, 65]. As a result, measuring this multiphoton transition using two-tone spectroscopy allows $E_{C}$ to be determined in situ for each qubit. For the three qubits, we find $E_{C}=2\pi\times 125$ MHz, $2\pi\times 113$ MHz, and $2\pi\times 122$ MHz, respectively.
The junctions that form the SQUID loop of the transmon are deposited via double-angle evaporation (see Appendix F for a detailed description of the fabrication process). As the junctions for the three transmons are fabricated simultaneously, the junctions must all share the same orientation. Thus, while the capacitor paddles are always parallel to the center pin of the coupled resonator, the junctions are instead parallel to the edge of the chip, leading to a minor difference in the designs for the on-axis qubit (Q2) and the off-axis qubits (Q1 and Q3). The transmons are designed to have tuning ranges that span from below the half-wave modes up to the full-wave modes. To experimentally determine $E_{J,sum}$ of the fabricated junctions, we match Eq. 21 and Eq. 23 to two-tone spectroscopy scans that map out the flux dependence of the transmon frequency and the photonic bound states (see Appendix C for details). For the three qubits, we find $E_{J,sum}=2\pi\times 100$ GHz, $2\pi\times 104$ GHz, and $2\pi\times 103$ GHz, respectively.
@@
Resonator $|t_{1}|$ $|t_{2}|$
Finite Element $2\pi\times 47$ MHz $2\pi\times 94$ MHz
DOS Comparison $2\pi\times 40$ MHz $2\pi\times 82$ MHz
Transmon $E_{C}$ $E_{J,sum}$
Finite Element $2\pi\times 165$ MHz -
Q1 (Measured) $2\pi\times 125$ MHz $2\pi\times 100$ GHz
Q2 (Measured) $2\pi\times 113$ MHz $2\pi\times$ 104GHz
Q3 (Measured) $2\pi\times 122$ MHz $2\pi\times 103$ GHz
Qubit-Resonator Coupling $g_{1}$ $g_{2}$
Finite Element $2\pi\times 76$ MHz $2\pi\times 152$ MHz
Measurement $2\pi\times 82.5$ MHz $2\pi\times 165$ MHz
Appendix C Determination of $g$
@@
Figure 11: Comparison of $g$ and the photonic bound states. (a) Transmission and two-tone qubit-spectroscopy data near the full-wave modes of the lattice. (Reproduced from Fig. 4) for Q3. The dashed line indicates the location of the bare-qubit transition used for the theoretical calculation in (b). (b) Theoretical calculation of the location of photonic bound states (red/orange) versus flux for a transmon (approximated as flux-tunable two-level system) coupled to the lattice with a coupling strength $g_{2}=2\pi\times 165$ MHz. The size and color of the marker indicates the fraction of the wave function which is qubit-like. Photon-like eigenstates not shown. The grey dashed line indicates the flux dependence of the bare qubit transition. The locations of the photonic bound states in (a) and (b) are in good agreement, indicating that qubits in our device have single-cavity coupling strengths $g_{2}\approx 2\pi\times 165$ MHz and $g_{1}=g_{2}/2\approx 2\pi\times 82.5$ MHz.
@@
In a conventional single-mode circuit QED system, the qubit-cavity coupling strength $g_{0}$ could be determined by measuring the vacuum Rabi splitting when the qubit frequency is swept through resonance with the mode [9]. Alternatively, the coupling could be determined from the dispersive shifts (or Stark shifts) on the qubit due to a detuned cavity. However, in our device, both of these methods can be problematic. For the resonant case, avoided crossings between the qubits and the normal modes of the lattice are clearly visible in Figs. 6, 7, 8, and 9. However, the lattice modes are close enough together in frequency that these avoided crossings are not independent. The Rabi split peaks from the qubit interacting with mode $k$ can undergo avoided crossings with those from mode $k+1$ and hide the true size of the splitting. These avoided crossings, therefore, cannot be directly used to reliably extract the normal mode couplings $g_{k,j}$. Additionally, the normal-mode couplings $g_{k,j}$ cannot be used to determine the single-cavity coupling $g_{0}$ without detailed knowledge of the normal-mode wave functions in the presence of disorder.
@@
For the far-detuned case, the photon-number dependent dispersive shift, or AC Stark shift, on qubit $j$ scales as $n*g_{k,j}^{2}$, where $n$ is the intracavity microwave photon number in mode $k$ [9]. Provided a strong enough qubit-cavity coupling, this shift produces resolvable peaks in the qubit transition frequency for each value of the intracavity photon number; the frequency difference between the transitions for $n$ photons and $n+1$ photons can subsequently be used to extract $g_{0}$ [67]. In our device, the normal-mode couplings involved are small enough that the Stark shifts on the qubits are not single-photon resolved and therefore cannot be used to calibrate $g_{k,j}$ without an independent measure of the intracavity photon number.
@@
Good qualitative agreement between the experimentally-measured and the numerically-simulated photonic bound states is found for $g_{2}=2\pi\times 165$ MHz at the full-wave modes. The corresponding coupling strength at the half-wave modes is then inferred to be $g_{1}=g_{2}/2=2\pi\times 82.5$ MHz. These values are close to the values of $g_{1}=2\pi\times 76$ MHz and $g_{2}=2\pi\times 152$ MHz derived from finite element simulations. The discrepancy is likely due to slight over or under-exposure of the design during fabrication.
A more quantitative technique to extract $g_{k,j}$ from non-linear spectroscopy is the subject of upcoming work in Ref. [68].
Appendix D Transmon Ionization and Mode-Mode Spectroscopy
In Section V.2, we presented the novel measurement technique mode-mode spectroscopy in which lattice modes can be detected with high SNR using their qubit-induced nonlinearity. In this method the response of a strongly driven mode is detected via the frequency shift induced on a weakly-driven monitor mode. At low drive powers, mode-mode spectroscopy signal originates from Kerr nonlinearity. However, the mode-mode spectroscopy presented in this work is collected at comparatively high drive powers, where the frequency shift of the monitor mode primarily arises from transmon ionization.
Transmon ionization is a process in which a transmon is excited above its cosine potential well. In conventional transmon readout, this process leads to measurement errors [62, 61]. When the drive amplitude applied to the resonator is high enough, the transmon can be excited directly to higher energy states through a multiphoton process, even with a far off-resonant drive. Here, the same principle is exploited to perform high-power mode-mode spectroscopy on our device. We apply a scanning drive tone which does not affect the transmon when off resonance, but when it comes into resonance with one of the photonic modes of the lattice, the resulting intracavity photon number is large enough to ionize one or more of the transmons. The resulting frequency shifts of the monitor mode, and associated transmission changes, are much larger than what is observed at low drive power, due to the lowest-order Kerr non-linearity alone.
Appendix E Flux calibration
In order to carry out all of the measurements above with the three qubits at well-known frequencies, a feed-forward scheme was implemented to remove cross-talk between the on-chip flux bias lines, following the procedure outlined in Ref. [69]. By first measuring and then inverting the cross-talk between the lines, a software correction can be applied in order to tune each qubit independently. This allows individual qubits to be tuned to well-determined detunings from the different photonic bands.

### v2

@@ Appendix B Supplementary Transmission Data
Figure 8: Low-power half-wave transmission data. (a)-(b) Transmission data of the half-wave modes as a function of flux applied to Q1 while the remaining two qubits are detuned from the range of interest: (a) transmission magnitude near the degenerate flat bands and lowest-energy dispersive band and (b) transmission magnitude over the full set of half-wave modes. (c)-(d) Analogous data for flux applied to Q2. (e)-(f) Analogous data for flux applied to Q3. The mutiphoton-transition and transmon-ionization glitches seen in Fig. 6 are not present here due to the lower input power used (55 dB lower for the zoomed-in scans, $10$ dB lower for the wide scans), and the avoided crossings are smooth and continuous. Due to the intra-unit-cell wavefunctions in the different bands and the inequivalent locations of the three qubits, Q2 does not couple to the half-wave flat bands, and as a result, the large avoided crossing visible for Q1 and Q3 in (a) and (e), is not present when tuning Q2 in (c). Additionally, the mode at $4.95$ GHz is an example of a mode which has a node in one unit cell, and therefore does not couple to Q1 in (b), but does couple to Q3 in (f), which is located on the equivalent lattice site one unit cell over.
@@
Since large probe powers can lead to non-linear effects, a tradeoff must be made between visibility of small features and clean avoided crossings. At high powers, modes are clearly visible with few averages and low-transmission modes can be detected with higher confidence. However, when a qubit is swept through resonance with a mode the drive tone has the potential to ionize the qubit (see Appendix F) leading to glitches in the avoided crossings. At low powers transmon ionization is less likely but achieving equivalent SNR for small features is generally not possible, or extremely slow. The crossing between the half-wave mode at $\sim 4.96$ GHz and Q3 is a perfect example of this effect of measurement-tone power. In a higher-power scan, such as Fig. 6(c), the avoided crossing "breaks" when Q3 is close to resonance with the mode, but in a scan taken with $10$ dB lower drive power (Fig. 8(f)), the full crossing is cleanly visible.
@@
Numerical simulations of the photonic bound states, shown in Fig. 11 and discussed in Appendix E, indicate that Q3, and by similarity Q1, attain peak frequencies of roughly 9.859.859.99.9 GHz, which are inside the full-wave band structure. As a result, modes in the highest dispersive band do not generally show full avoided crossings, merely deflections. Modes in this upper dispersive band are particularly difficult to observe directly in transmission, due in part to the broad parasitic modes situated within the band. For these modes, the phase response of the transmitted signal shown in Fig. 9(a)-(c)(iii) is stronger than the amplitude response that is typically used to characterize the other photonic modes.
Appendix C Two-Tone Spectroscopy
The dispersive readout mechanism described in Section III.1 can be used as a spectroscopy technique for locating qubit transitions by combining two drive tones: one to excite the qubit and one to monitor the qubit state. In the case of conventional single-mode devices, the fixed-frequency monitor tone is applied on-resonance with a resonator, or a normal mode in the case of the CPW lattice device studied here, whereas the drive/excitation tone is swept in frequency. This technique, known simply as two-tone spectroscopy, is standard in the field [Blais:revmodphys], and is conceptually simple in the idealized case in which the drive is known to implement a single $\pi$ pulse and state readout using the monitor tone is done in a time-resolved fashion. The base temperature of the dilution refrigerator naturally prepares the qubit in its ground state, and if the drive tone is resonant with the qubit transition, then the qubit will be excited, causing the homodyne signal at the monitor frequency to switch to a different amplitude and phase.
[…]
Appendix D General Device Design
[…]
D.1 On-Site Energies
[…]
@@ D.1 On-Site Energies
Another potential source of disorder arises from the capacitance between neighboring resonators, since the coupling capacitors on the device, highlighted blue in Fig. 10, induce on-site energy shifts in the resonators. To keep this contribution uniform, all resonators must have the same number of coupling capacitors [Koch:TRS_breaking, Kollar:2019hyperbolic], two at each end for the case of the device shown here. Therefore, the same 3-way coupling-capacitor geometry is used at the ends of all resonators, both for sites in the bulk which naturally have four neighbors, as well as for sites at the edge of the device, where some of the neighboring resonators are absent. Unused ports on the 3-way capacitors are either connected to input/output feedlines, or terminated with a high-frequency $\lambda/4$ resonator, which does not participate in the lattice due to the large detuning.
@@
D.2 Hopping Parameters
To determine the predicted zeroth-order hopping rate in our device, we use the resonant frequencies determined from the shape-test devices (or finite element simulations in HFSS), $Z_{0}=50\ \Omega$, and a finite element simulation of the coupling capacitance using Ansys Maxwell. Our three-way coupler with a gap of $2\penalty 10000\ \mu m$ is simulated to have a $19.5$ fF capacitance between each pair of resonators, from which $|t_{1}^{(0)}|/2\pi$ is predicted to be $\sim 47$ MHz and $|t_{2}^{(0)}|/2\pi$ is predicted to be $\sim 94$ MHz. For the main device, the empirically determined hopping rates for the half-wave and full-wave modes are determined through comparison of the experimentally-measured photonic bands with the frequency-dependent hopping model of the target lattice (see Section V.2). These values of $|t_{1}^{(0)}|/2\pi\approx 40$ MHz and $|t_{2}^{(0)}|/2\pi\approx 82$ MHz are reasonably close to the simulated values predicted by circuit theory and finite-element simulation. The discrepancy may be due to slight over-exposure of the device during photolithography, or due to inaccuracy in the Ansys Maxwell capacitance simulation at the level of $\sim 1$ fF.
D.3 Transmon Design
Resonator $|t_{1}|/2\pi$ $|t_{2}|/2\pi$
Finite Element $47$ MHz $94$ MHz
DOS Comparison $40$ MHz $82$ MHz
Transmon $E_{C}/2\pi$ $E_{J,sum}/2\pi$
Finite Element $165$ MHz -
Q1 (Measured) $125$ MHz $100$ GHz
Q2 (Measured) $113$ MHz $104$ GHz
Q3 (Measured) $122$ MHz $103$ GHz
Qubit-Resonator Coupling $g_{1}/2\pi$ $g_{2}/2\pi$
Finite Element $76$ MHz $152$ MHz
Measurement $82.5$ MHz $165$ MHz
Before the design of the transmon qubits embedded in the CPW lattice is discussed in detail, we first introduce some of the traditional parameters of a transmon, as defined in Refs. [Koch:transmon, Blais:revmodphys, quantumengineersguide]. The frequency of the transition between the ground and first-excited state is approximately given by
@@ D.3 Transmon Design
The total capacitance takes into account not only the direct capacitance between the two paddles but also the indirect coupling through the center pin of the CPW resonator and the ground planes of the device [quantumengineersguide]. For flux-tunable transmons, the effective total $E_{J}$ is dependent on the individual Josephson energies of the two junctions in the SQUID loop and the external applied flux. The full form for the effective $E_{J}$ of a tunable transmon is given by
@@
In order to suppress the effect of charge noise in the substrate, a transmon must operate in the regime where $E_{J}/E_{C}\gg 1$, typically $20$ or larger [Koch:transmon, Blais:revmodphys]. This requires a large $C_{\Sigma}$, which can either be achieved using large-area capacitor paddles or a digitated capacitor. Our design uses a digitated capacitor in order to make the qubit compact and easier to incorporate in different resonator geometries. Conventional, non-digitated designs would need to be significantly wider, as much as 5105-10 times, to achieve an equivalent capacitance, which would severely restrict the ability to tightly pack resonators in the array [Gambetta:fab]. For our design, the total capacitance between the paddles is determined to be $C_{\Sigma}=0.1174$ pF using Ansys Maxwell. Following from Eq. 23, $E_{C}/2\pi$ is predicted to be $165$ MHz. In a transmon, the two-photon transition from the ground state to the second-excited state is lower in frequency than the main transition by $\sim E_{C}/2$ [Blais:revmodphys, quantumengineersguide]. As a result, measuring this multiphoton transition using two-tone spectroscopy allows $E_{C}$ to be determined in situ for each qubit. For the three qubits, we find $E_{C}/2\pi=125$ MHz, $113$ MHz, and $122$ MHz, respectively.
The junctions that form the SQUID loop of the transmon are deposited via double-angle evaporation (see Appendix H for a detailed description of the fabrication process). As the junctions for the three transmons are fabricated simultaneously, the junctions must all share the same orientation. Thus, while the capacitor paddles are always parallel to the center pin of the coupled resonator, the junctions are instead parallel to the edge of the chip, leading to a minor difference in the designs for the on-axis qubit (Q2) and the off-axis qubits (Q1 and Q3). The transmons are designed to have tuning ranges that span from below the half-wave modes up to the full-wave modes. To experimentally determine $E_{J,sum}$ of the fabricated junctions, we match Eq. 22 and Eq. 24 to two-tone spectroscopy scans that map out the flux dependence of the transmon frequency and the photonic bound states (see Appendix E for details). For the three qubits, we find $E_{J,sum}/2\pi=100$ GHz, $104$ GHz, and $103$ GHz, respectively.
@@
Appendix E Determination of $g$
@@
Figure 11: Comparison of $g$ and the photonic bound states. (a) Transmission and two-tone qubit-spectroscopy data near the full-wave modes of the lattice. (Reproduced from Fig. 4) for Q3. The dashed line indicates the location of the bare-qubit transition used for the theoretical calculation in (b). (b) Theoretical calculation of the location of photonic bound states (red/orange) versus flux for a transmon (approximated as flux-tunable two-level system) coupled to the lattice with a coupling strength $g_{2}/2\pi=165$ MHz. The size and color of the marker indicates the fraction of the wave function which is qubit-like. Photon-like eigenstates not shown. The grey dashed line indicates the flux dependence of the bare qubit transition. The locations of the photonic bound states in (a) and (b) are in good agreement, indicating that qubits in our device have single-cavity coupling strengths $g_{2}/2\pi\approx 165$ MHz and $g_{1}/2\pi=(g_{2}/2)/2\pi\approx 82.5$ MHz.
@@
In a conventional single-mode circuit QED system, the qubit-cavity coupling strength $g_{0}$ could be determined by measuring the vacuum Rabi splitting when the qubit frequency is swept through resonance with the mode [Blais:revmodphys]. Alternatively, the coupling could be determined from the dispersive shifts (or Stark shifts) on the qubit due to a detuned cavity. However, in our device, both of these methods can be problematic. For the resonant case, avoided crossings between the qubits and the normal modes of the lattice are clearly visible in Figs. 6, 7, 8, and 9. However, the lattice modes are close enough together in frequency that these avoided crossings are not independent. The Rabi split peaks from the qubit interacting with mode $k$ can undergo avoided crossings with those from mode $k+1$ and hide the true size of the splitting. These avoided crossings, therefore, cannot be directly used to reliably extract the normal mode couplings $g_{k,j}$. Additionally, the normal-mode couplings $g_{k,j}$ cannot be used to determine the single-cavity coupling $g_{0}$ without detailed knowledge of the normal-mode wave functions in the presence of disorder.
@@
For the far-detuned case, the photon-number dependent dispersive shift, or AC Stark shift, on qubit $j$ scales as $n*g_{k,j}^{2}$, where $n$ is the intracavity microwave photon number in mode $k$ [Blais:revmodphys]. Provided a strong enough qubit-cavity coupling, this shift produces resolvable peaks in the qubit transition frequency for each value of the intracavity photon number; the frequency difference between the transitions for $n$ photons and $n+1$ photons can subsequently be used to extract $g_{0}$ [Schuster_2007:Number_Splitting]. In our device, the normal-mode couplings involved are small enough that the Stark shifts on the qubits are not single-photon resolved and therefore cannot be used to calibrate $g_{k,j}$ without an independent measure of the intracavity photon number.
@@
Good qualitative agreement between the experimentally-measured and the numerically-simulated photonic bound states is found for $g_{2}/2\pi=165$ MHz at the full-wave modes. The corresponding coupling strength at the half-wave modes is then inferred to be $g_{1}/2\pi=(g_{2}/2)/2\pi=82.5$ MHz. These values are close to the values of $g_{1}/2\pi=76$ MHz and $g_{2}/2\pi=152$ MHz derived from finite element simulations. The discrepancy is likely due to slight over or under-exposure of the design during fabrication.
A more quantitative technique to extract $g_{k,j}$ from non-linear Kerr spectroscopy is the subject of upcoming work in Ref. [KerrSpec].
Appendix F Transmon Ionization and Mode-Mode Spectroscopy
In Section V.2, we introduced the measurement technique mode-mode spectroscopy in which lattice modes can be detected with high SNR using their qubit-induced nonlinearity. In this method the response of a strongly driven mode is detected via the frequency shift induced on a weakly-driven monitor mode. At low drive powers, the signal originates from Kerr nonlinearity [Peugeot_2024, Bosman_2017]. However, the mode-mode spectroscopy presented in this work is collected at comparatively high drive powers, where the frequency shift of the monitor mode primarily arises from transmon ionization.
Transmon ionization is a process in which a transmon is excited above its cosine potential well. In conventional transmon readout, this process leads to measurement errors [Blais:ionization, Martinis:ionization]. When the drive amplitude applied to the resonator is high enough, the transmon can be excited directly to higher energy states through a multiphoton process, even with a far off-resonant drive. Here, the same principle is exploited to perform high-power mode-mode spectroscopy on our device. We apply a scanning drive tone which does not affect the transmon when off resonance, but when it comes into resonance with one of the photonic modes of the lattice, the resulting intracavity photon number is large enough to ionize one or more of the transmons. The resulting frequency shifts of the monitor mode, and associated transmission changes, are much larger than what is observed at low drive power, due to the lowest-order Kerr non-linearity alone.
Appendix G Flux calibration
In order to carry out all of the measurements above with the three qubits at well-known frequencies, a feed-forward scheme was implemented to remove cross-talk between the on-chip flux bias lines, following the procedure outlined in Ref. [Ma:2019ba]. By first measuring and then inverting the cross-talk between the lines, a software correction can be applied in order to tune each qubit independently. This allows individual qubits to be tuned to well-determined detunings from the different photonic bands.